OpenStudy (anonymous):

If the series of an and the series of b are positive term series for which series an converges and series bn diverges. Would series an*bn converge, diverge, or does it have not sufficient information?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

lets try a couple examples

OpenStudy (misty1212):

\[\sum\frac{1}{n^3}\] converges right ?

OpenStudy (misty1212):

this is where you say "yeah it does"

OpenStudy (anonymous):

Yup!

OpenStudy (misty1212):

and how about \[\sum n\]

OpenStudy (anonymous):

diverges, it does to infinity

OpenStudy (misty1212):

multiply and get \[\sum\frac{1}{n^2}\] converge or diverge?

OpenStudy (anonymous):

converge

OpenStudy (anonymous):

But that is only one example....

OpenStudy (misty1212):

right but now at least we know it CAN converge now try it with \[\sum\frac{1}{n^2} , \sum n^2\]

OpenStudy (anonymous):

converge once again!

OpenStudy (misty1212):

oh no!

OpenStudy (misty1212):

\[\frac{1}{n^2}\times n^2=1\] and \[\sum 1\] does not converge

OpenStudy (anonymous):

Wait I mean diverge since it is a positive term

OpenStudy (anonymous):

Divergence test shows that it goes to infinity.

OpenStudy (anonymous):

Okay so we showed it could be divergent of convergent.... so this is not enough information?

OpenStudy (anonymous):

Okay, thank you!

OpenStudy (misty1212):

yeah it totally depends on each \[\color\magenta\heartsuit\]